# Entropic signature of resonant thermal transport: Ordered form of heat conduction

### v3

@@ V Heat flux autocorrelation in the presence of resonances
Figure 4: Heat flux ACF in NPMs. The heat flux is extracted from equilibrium simulations at room temperature for (a) the membrane and (b-e) the NPM with different nanopillar heights. We extract the heat flux from the entire simulation domains (b,d) and only from the base-membrane portion (c,e).
@@
The strong current fluctuations in NPMs are associated with resonant hybridizations of the propagating phonon modes in the base membrane. In this section, we explore the origin and manifestation of this phenomenon using equilibrium MD simulations. Specifically, we first thermalize the NPM system following the protocol described in Appendix C. Thereafter and under the resulting equilibrium conditions at room temperature, we analyze the heat flux autocorrelation function (ACF) $<\textbf{J}(0)\otimes\textbf{J}(t)>$, where J is the heat flux vector integrated over the volume and sampled every 4 fs and $t$ is the delay time. For each nanostructure, we perform the simulations in the supercell consisting of five unit cells also considered in the previous section.
@@
In Fig. 4, we show the heat flux ACF for different nanopillar heights $h$. First, we investigate the evolution of the global heat flux by integrating J over the volume of the entire NPM system. We observe that the autocorrelation of the thermal fluctuations decays completely over a time scale of 10s of picoseconds, which indicates that the atom dynamics in different regions of the system are not fully correlated. This suggests that thermodynamic equilibrium with ergodic evolution is robustly established in our simulations. Even though our simulation domains are relatively small and sustain strong dynamical constraints, the mixing condition is thus satisfied, which in turn justifies the applicability of the fluctuation theorem and the TUR.
@@
To examine the influence of resonances in more detail, in Fig. 4 we also show the ACF by extracting the heat flux J only from the base-membrane portion of the NPM. The local heat flux ACF displays a remarkably slow initial decay followed by periodic oscillations that persist over a nanosecond time scale. This phenomenon is not observed in the bulk, membrane, or NPC systems, and becomes more pronounced in the presence of taller nanopillars, where the resonant effect is enhanced. Therefore, these features are a manifestation of coherent phonon dynamics in NPMs, and further confirm that the energy transport mechanisms in the base membrane are strongly influenced by the phonon coupling with the resonant modes in the nanopillars. In analogy with these results, previous work demonstrated that similar features in the heat flux ACF can appear in core-shell nanowires due to the emergence of resonant transversal modes that couple with the propagating longitudinal modes [54]--a mechanism that is distinctly different from the resonance hybridizations that occur in NPMs, yet still brings rise to this type of behavior.
@@
Therefore, the anomalous behavior of the heat flux ACF under the influence of atomic-scale phonon resonances at equilibrium is a counterpart of the enhancement of the energy transfer fluctuations in non-equilibrium conditions (see Fig. 3). The coherent coupling of the phonon modes in the nanopillars and the base membrane impairs the stochastic evolution of the vibrational energy distribution, which hinders the emergence of thermodynamic equilibrium. This suggests that the ergodic exploration of the phase space is not as easily accessible in material systems presenting resonances compared to conventional materials, and might indicate a physical origin underlying errors and uncertainty in Green-Kubo calculations [55, 56] to extract the thermal conductivity in complex nanoscale configurations prone to various forms of resonances [54]. Nevertheless, quantification of the heat flux uncorrelation time and the conductivity was shown possible in NPMs with certain dimensions [57, 58] and other complex systems [59].
@@ VI Entropy maximization time
This metric characterizes the time required by a system to evolve from the minimum-entropy state, corresponding to a fully coherent and non-thermalized state, towards equilibrium at the given temperature. However, $\tau_{S}$ does not correspond to actual thermalization times because the production rates $\widehat{\Sigma}_{\tau}/\tau$ are characterized in steady-state conditions at 300 K, whereas the entropy evolution is a function of temperature and sensitive to the specific non-equilibrium transient trajectory. Nevertheless, this idealized time enables equal-footing comparison of the influence of different dynamical constraints on the emergence of thermal equilibrium.
@@ Conclusion
In conclusion, we introduced a novel framework for the characterization of nanoscale heat conduction phenomena based on quantification of the entropic fingerprint of phonon motion, and we demonstrated the general concept of reversible thermal transport in the presence of nanoscale dynamical constraints. Primarily, the analysis reveals that the inclusion of intrinsically resonating nanostructures enables a regime of ordered heat conduction to be approached--one that is characterized by a strongly fluctuating energy current, where the randomization and thermalization of the system--rigorously quantified by the rate of entropy production--are hindered and the coherent response of the phonons is distinctly prevalent. Leveraging these effects opens a new horizon for nanoscale phonon engineering with potential implications on a wide range of applications including semiconductor technology, thermoelectricity, superconductivity [60], and emerging quantum devices [61]--all that would benefit from controlled phonon evolution.
Figure 5: Entropy maximization time in the different nanosystems assuming a constant entropy production rate.
@@ Appendix C Equilibrium MD and configurational entropy
We perform equilibrium MD simulations to characterize the configurational entropy distributions in the different nanosystems. A single unit cell is simulated at 300 K. The systems are initially equilibrated for 1 ns using a time step 0.5 fs under the NPT ensemble, thus allowing the reconfiguration of the surfaces due to boundary atoms relaxation. The excess entropy of each atom is sampled every 6 ps and averaged during the subsequent 6 ns under the NVE ensemble. As discussed in [27], the excess entropy of atom $i$ reads
@@ C.2 Internal energy and configurational entropy
Appendix D Non-equilibrium MD and entropy production
We perform non-equilibrium MD simulations to compute entropy production lower bounds in each nanosystem. A supercell consisting of five aligned unit cells is simulated in steady state. To prevent rotational or translational effects, a 1.09 nm-thick layer of fixed silicon atoms are added to two opposite ends of each system. Using a Langevin thermostat and a time step of 0.5 fs, the rest of the system is first thermalized at 300 K for 1 ns. Then, a thermal gradient is imposed by thermalizing 2.17 nm-thick layers of atoms in between the fixed ends and the supercell using a Langevin thermostat at 310 K (source) and 290 K (sink), respectively, while leaving the supercell atoms free. The steady state, characterized by a stationary temperature profile between source and sink, is reached by running the simulation for 1 ns with a time step 0.5 fs. Thereafter, stationary evolution is simulated using the same time step during the subsequent 30 ns, wherein the energy transfer between thermostats is evaluated. Specifically, the energy transferred $q_{\tau}$ corresponds to the total energy introduced in the source to maintain the system out-of-equilibrium in steady state during a time window of $\tau=1$ ns. The values of $\text{E}(q_{\tau})$ and $\text{Var}(q_{\tau})$ required to calculate the entropy production bound $\widehat{\Sigma}_{\tau}$ are the mean and the variance of the energy $q_{\tau}$, respectively. The results are averaged over five simulations with different initial conditions to extract these statistical moments in each case. Similar results are obtained if considering the energy removed from the sink instead of the energy injected in the source to define $q_{\tau}$.
@@ D.1 Size effects
Due to the small size of the simulation domain, incorrect sampling of low frequency phonons, nonphysical coupling between thermostats, and enhanced ballistic phonon transmission can unfold. This would influence the energy transmission within the system [63, 64], and thus might impact the entropy production characterization. In Fig. 9, we show $\widehat{\Sigma}_{\tau}$ as a function of the number of unit cells $D$ between thermostats in the membrane geometry. After adequate normalization by the variable average thermal gradient as per Eq. (4), we do not observe significant size-dependent entropy production per atom in the membrane for $D>$ 24 CCs. This indicates that the results considering five unit cells between thermostats (i.e., $D=30$ CCs), are not strongly influenced by the small size of the simulation domain. Moreover, we consider the same distance between thermostats for all the nanostructures, and hence the results illustrated in Fig. 3 provide an equal-footing comparison among the different nanostructures. We note that $\widehat{\Sigma}_{\tau}$ is converged for relatively small system sizes $D$ relative to the direct calculation of the thermal conductivity using non-equilibrium simulations [64]. We attribute this to the functional form of the entropy bound, where the energy current $q_{\tau}$ is normalized by combining the mean and the variance. While this significantly reduces the computational cost of the $\widehat{\Sigma}_{\tau}$ comparison between nanostructures, the present simulation setting might be inadequate to directly compute converged values of the thermal conductivity using the average heat current.
@@ D.2 Influence of temperature difference between thermostats
Figure 12: Entropy production normalized by the number of atoms. Calculations are done over a time-window $\tau$=1 ns.
@@ D.4 Normalized entropy production
The entropy maximization and thermalization in our nanosystem models are achieved via anharmonic phonon-phonon interactions. Hence, the entropy production of a nanosystem is expected to decrease by reducing its size to scales smaller than the phonon mean free paths and scattering times. In fact, we obtain very strong reductions in the production of entropy in the nanostructured geometries relative to the bulk. This trend explains why, in general, non-equilibrium thermodynamic behavior tends to persist in semiconductors at reduced length or time scales.

### v4

@@ V Heat flux autocorrelation in the presence of resonances
The strong current fluctuations in NPMs are associated with resonant hybridizations of propagating phonon modes in the base membrane. In this section, we explore the origin and manifestation of this phenomenon using equilibrium MD simulations. Specifically, we first thermalize the systems following the protocol described in Appendix C. Thereafter and under the resulting equilibrium conditions at room temperature, we analyze the heat flux autocorrelation function (ACF) $<\textbf{J}(0)\otimes\textbf{J}(t)>$, where $t$ is the delay time and J is the heat flux vector integrated over the volume and sampled every 4 fs. For each nanostructure, we perform the simulations in the supercell consisting of five unit cells also considered in the previous section.
Figure 4: Heat flux ACF. The heat flux is extracted from equilibrium simulations at room temperature for (a) the bulk, (b) the membrane and (c-f) the NPM with different nanopillar heights. The heat flux is obtained from the entire simulation domains (c,e) and only from the base-membrane portion (d,f). Direct comparisons between the various ACF functions are illustrated in (g,h). The flux ACF in the base membrane of an NPM displays remarkably slow decay followed by an oscillatory behavior, which indicates the strong coherent coupling between the membrane and the nanopillar.
@@
In Fig. 4, we show the heat flux ACF for the bulk material, the membrane, and two cases of NPMs with different nanopillar heights $h$. We observe that the autocorrelation of the thermal fluctuations in the NPMs decays faster than in the bulk or the membrane cases, and tends to zero over a time scale of tens of picoseconds. This indicates that the atom dynamics in different regions of the NPM system are not fully correlated, and suggests that thermodynamic equilibrium with ergodic evolution is robustly established in our simulations. Even though our simulation domains are relatively small and sustain strong dynamical constraints, the mixing condition is thus satisfied, which in turn justifies the applicability of the fluctuation theorem and the TUR.
@@
To examine the influence of resonances in more detail, in Fig. 4 we also show the ACF by extracting the heat flux J only from the base-membrane portion of the NPM. The local heat flux ACF displays a remarkably slow initial decay followed by periodic oscillations that persist over a nanosecond time scale. This phenomenon is not observed in the bulk, membrane, or NPC systems, and becomes more pronounced in the presence of taller nanopillars, where the resonant effect is enhanced. Therefore, these features are a manifestation of coherent phonon dynamics in NPMs, and further confirm that the energy transport mechanisms in the base membrane are strongly influenced by the phonon coupling with the resonant modes in the nanopillars. More specifically, the heat flux ACF in the base membrane of an NPM decays slower than in the membrane case. This indicates that the NPM resonances, in fact, work towards increasing the phonon correlations, rather than reducing them as done by scattering from non-pristine boundaries. In analogy with these results, previous work demonstrated that similar features in the heat flux ACF can appear in core-shell nanowires due to the emergence of resonant transversal modes that couple with the propagating longitudinal modes [59]--a mechanism that is distinctly different from the resonance hybridizations that occur in NPMs, yet still brings rise to this type of behavior.
@@
The anomalous behavior of the heat flux ACF under the influence of atomic-scale phonon resonances at equilibrium is a counterpart of the enhancement of the energy transfer fluctuations in nonequilibrium conditions (see Fig. 3). The coherent coupling of the phonon modes in the nanopillars and the base membrane impairs the stochastic evolution of the vibrational energy distribution, which obstructs the emergence of thermodynamic equilibrium. This suggests that ergodic exploration of the phase space is not as easily accessible in material systems presenting resonances compared to conventional materials, and might indicate a physical origin underlying errors and uncertainty in Green-Kubo calculations [60, 61] to extract the thermal conductivity in complex nanoscale configurations prone to various forms of resonances [59]. Nevertheless, quantification of the heat flux uncorrelation time and the conductivity was shown possible in NPMs with certain dimensions [62, 63] and other complex systems [64].
@@ VI Entropy maximization time
This metric characterizes the time required by a system to evolve from the minimum-entropy state, corresponding to a fully coherent and non-thermalized state, towards equilibrium at the given temperature. However, $\tau_{S}$ does not correspond to actual thermalization times because the production rates $\widehat{\Sigma}_{\tau}/\tau$ are characterized in steady-state conditions at 300 K, whereas the entropy evolution is a function of temperature and sensitive to the specific nonequilibrium transient trajectory. Nevertheless, this idealized time enables equal-footing comparison of the influence of different dynamical constraints on the emergence of thermal equilibrium.
Figure 5: Entropy maximization time in the different nanosystems assuming a constant entropy production rate. The larger configurational entropy and reduced entropy production rate obtained in NPMs combine to slow down the thermalization process with respect to the other material systems.
@@ Conclusion
In conclusion, we introduced a novel framework for the characterization of nanoscale heat conduction phenomena based on quantification of the entropic fingerprint of phonon motion. By equilibrium simulations, we evaluated the influence of nanoscale dynamical constraints on the available phase space for atomic motion, highlighting the additional configurational entropy required to thermalize different nanostructure features when compared to the bulk form. By nonequilibrium simulations, we demonstrated that dynamical constraints induce a degree of reversible evolution in the presence of nanoscale energy currents. Primarily, our findings reveal that the inclusion of intrinsically resonating nanostructures enables a regime of ordered heat conduction to be approached. This regime is characterized by a strongly fluctuating energy current, where the randomization and thermalization of the system--rigorously quantified by the rate of entropy production--are hindered and the coherent response of the phonons is distinctly prevalent. Leveraging these effects opens a new horizon for nanoscale phonon engineering with potential implications on a wide range of applications including semiconductor technology, thermoelectricity [65], superconductivity [66], and emerging quantum devices [67]--all that would benefit from controlled phonon evolution.
@@ Appendix C Equilibrium MD and configurational entropy
We perform equilibrium MD simulations at 300 K to characterize the configurational entropy distributions and the heat flux ACF in the different nanosystems. A single unit cell is simulated for the configurational entropy calculations, and a supercell consisting of five unit cells is considered for the heat flux ACF calculations. The systems are initially equilibrated for 1 ns using a time step 0.5 fs under the NPT ensemble, thus allowing the reconfiguration of the surfaces due to boundary atoms relaxation.
The excess entropy of each atom is sampled every 6 ps and averaged during the subsequent 6 ns under the NVE ensemble. As discussed in [26], the excess entropy of atom $i$ reads
@@ C.2 Internal energy and configurational entropy
Appendix D Nonequilibrium MD and entropy production
We perform nonequilibrium MD simulations to compute entropy production lower bounds in each nanosystem. A supercell consisting of five aligned unit cells is simulated in steady state. To prevent rotational or translational effects, a 1.09 nm-thick layer of fixed silicon atoms are added to two opposite ends of each system. Using a Langevin thermostat and a time step of 0.5 fs, the rest of the system is first thermalized at 300 K for 1 ns. Then, a thermal gradient is imposed by thermalizing 2.17 nm-thick layers of atoms in between the fixed ends and the supercell using a Langevin thermostat at 310 K (source) and 290 K (sink), respectively, while leaving the supercell atoms free. The steady state, characterized by a stationary temperature profile between source and sink, is reached by running the simulation for 1 ns with a time step 0.5 fs. Thereafter, stationary evolution is simulated using the same time step during the subsequent 30 ns, wherein the energy transfer between thermostats is evaluated. Specifically, the energy transferred $q_{\tau}$ corresponds to the total energy introduced in the source to maintain the system out-of-equilibrium in steady state during a time window of $\tau=1$ ns. The values of $\text{E}(q_{\tau})$ and $\text{Var}(q_{\tau})$ required to calculate the entropy production bound $\widehat{\Sigma}_{\tau}$ are the mean and the variance of the energy $q_{\tau}$, respectively. The results are averaged over five simulations with different initial conditions to extract these statistical moments in each case. Similar results are obtained if considering the energy removed from the sink instead of the energy injected in the source to define $q_{\tau}$.
@@ D.1 Size effects
Due to the small size of the simulation domain, incorrect sampling of low frequency phonons, nonphysical coupling between thermostats, and enhanced ballistic phonon transmission can unfold. This would influence the energy transmission within the system [69, 70], and thus might impact the entropy production characterization. In Fig. 9, we show $\widehat{\Sigma}_{\tau}$ as a function of the number of unit cells $D$ between thermostats in the membrane geometry. After adequate normalization by the variable average thermal gradient as per Eq. (4), we do not observe significant size-dependent entropy production per atom in the membrane for $D>$ 24 CCs. This indicates that the results considering five unit cells between thermostats (i.e., $D=30$ CCs), are not strongly influenced by the small size of the simulation domain. Moreover, we consider the same distance between thermostats for all the nanostructures, and hence the results illustrated in Fig. 3 provide an equal-footing comparison among the different nanostructures. We note that $\widehat{\Sigma}_{\tau}$ is converged for relatively small system sizes $D$ relative to the direct calculation of the thermal conductivity using nonequilibrium simulations [70]. We attribute this to the functional form of the entropy bound, where the energy current $q_{\tau}$ is normalized by combining the mean and the variance. While this significantly reduces the computational cost of the $\widehat{\Sigma}_{\tau}$ comparison between nanostructures, the present simulation setting might be inadequate to directly compute converged values of the thermal conductivity using the average heat current.
@@
Figure 12: Entropy production normalized by the number of atoms. As a reference, the inset shows the total entropy production.
@@ D.4 Normalized entropy production
The entropy maximization and thermalization in our nanosystem models are achieved via anharmonic phonon-phonon interactions. Hence, the entropy production of a nanosystem is expected to decrease by reducing its size to scales smaller than the phonon mean free paths and scattering times. In fact, we obtain very strong reductions in the production of entropy in the nanostructured geometries relative to the bulk. This trend explains why, in general, nonequilibrium thermodynamic behavior tends to persist in semiconductors at reduced length or time scales.